an ideal gas undergoes a single stage expansion against a constant opposing pressure from (P1,V1,T) to(P2,V2,T).what is the largest mass m which can be lifted through a height h in the expansion?the answer does nat have p ext ,so i am confused.

Grade:

4 Answers

Chetan Mandayam Nayakar
11 years ago

work done=nRTln(V2/V1)=mgh

anushka shanbhag
11 years ago

the answer is wrong.external pressure is constant so u cannot apply reversible process equation.

Sarthak Shrivastava
6 years ago
askIITians your IITians have provided a wrong answer. The external pressure will be P2. Hope with this you will be able to solve with this information
 
Aastha
2 years ago
Work done = mgh=Pext(nRT/P2-nRT/P1)
Given:expagainst constt.pressure.....irreversible process
Pext=Pfinal =P2
So, m=nRT/gh(1-P2/P1)